The risk-free rate of return is 5%, the expected rate of return on the market portfolio is 16%, and the stock of Xyrong Corporation has a beta coefficient of 1.4. Xyrong pa Dividends were just paid and are expected to be paid annually. You expect that Xyrong will earn an ROE of 25% per year on al reinvested earnings forever. ys out 60% of its earnings in dividends, and the latest earnings announced were $700 per share. a. What is the intrinsic value of a share of Xyrong stock? (Do not round intermediate calculetions. Round your answer to 2 decimal places.) Intrinsic value b. If the market price of a share is currently $40, and you expect the market price to be equal to the intrinsic value one year from now, what is your expected 1-year holding-period return on Xyrong stock? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Expected one-year holding-period return
Answer #1

1.

Growth rate=25%*(1-60%)=10%

Required return=5%+1.4*(16%-5%)=20.4%

Intrinsic value=7*60%*(1+10%)/(20.4%-10%)=44.42

2.

=(44.42*1.1+7*60%*(1+10%))/40-1=33.71%